OpenStudy (anonymous):

rewrite this using associative property, (w+8)(-5)

OpenStudy (anonymous):

er... you cant i see w plus 8 all multiplied by -5

OpenStudy (anonymous):

associative property of addition (a+b) +c = a+(b+c)

OpenStudy (anonymous):

Associative Property of Addition a + (b + c) = (a + b) + c EX.) 2 + (3 + 4) = (2 + 3) + 4

OpenStudy (anonymous):

Associative Property of Multiplication a(bc) = (ab)c EX.) 2(3×4) = (2×3)4

OpenStudy (anonymous):

-5(w+8) or (w+8)-5